OpenStudy (anonymous):

Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 7x - 14 = 6x + 2 Equation #2 7 - 2x = 5x + 21

OpenStudy (anonymous):

The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describe solutions to equations shown.

OpenStudy (anonymous):

7x - 6x -14 = 2 7x-6x = 14 +2 x = 16 7-2x - 5x =21 -2x -5x = 21 -7 -7x = 14 x=-2 The first solution is larger, so (b)

OpenStudy (anonymous):

does the whole process make sense to you? You can be honest